# Scaling to Multimodal and Multichannel Heart Sound Classification: Fine-Tuning Wav2Vec 2.0 with Synthetic and Augmented Biosignals

### v2

@@ 3.4.3 Multichannel models
The multichannel architecture is shown in Figure 13. The multichannel model includes a support vector machine (SVM) that replaces the hidden layers and output layer of the MLP after training has been complete, leaving only the initial fully connected layer of the MLP. Additionally, for fine-tuning, low-rank adaptation (LoRA) [35] is utilised for the multichannel model due to this dataset containing more limited data than the other datasets. Both the combination of LoRA for fine-tuning and the SVM help to prevent the model from overfitting to the training set.
@@ 4.1 Single channel models
The PaCMAP [36] plots of an average-performing model from each experiment are shown below to represent how well, on average, those models encode the data before going through the classifier. The data used for these plots is also from the test set, so these models have not been trained on this data. Figure 17, Figure 18, and Figure 19 illustrate the PaCMAP embeddings for the CinC models, with input data sampled at 16kHz with no augmented dataset, 16kHz with augmented dataset and 4.125kHz with augmented dataset, respectively.
@@ 4.2 Multimodal models
Comparing our model that utilises both PCG and ECG signals to other models that were evaluated on training-a, We can see that our model has the greatest performance with the highest for each metric, other than TPR, in Table 16. Although it does not achieve the highest TPR, our proposed model achieves a much higher TNR, than the CNN model from [5], resulting in all other metrics being greater and hence a better performing model. It is observed that the majority of methods within this comparison utilise the raw signal, as opposed to spectrograms, which is more common for models trained on the entire CinC dataset. However, none of the previous approaches utilised transformer architectures. As training-a is a small dataset, and hence, models trained on it will not generalise well, the results from training on the CinC dataset demonstrate that the method can effectively scale to larger amounts of data while still delivering performance improvements.
@@
CNN [5] Spectrogram 91.25% 84.17% 98.33% 70.00% 94.40%
RNN [42] Raw Signal 91.60% 91.55% 92.00% 91.10% 91.50%
@@ 4.3 Multichannel models
Comparing with the original study utilising this wearable vest, it is found that the performance is slightly reduced, however, this is not on the same dataset. The other method [22] made use of 10s data, where the patients held their breath whilst also relying on hand-annotated labels for the heart cycles. This method made use of subjects from an additional round of data collection in a different hospital ward while the subjects were breathing. This resulted in noisier data and hence a distribution shift compared to the first data collection round. Hence, it is found that this study is more practical, making use of easier-to-collect data and getting comparable performance on noisier data. In general, this proposed method is more flexible as it does not require manual heart cycle segmentation and does not require subjects to hold their breath, despite the slightly lower performance in some metrics seen in Table 19.
@@
SVM [22] 80 subjects (40CAD/40NOR), 10s breath held LFCCs 80.44% 80.435% 85.25% 75.62% 81.00%
@@
The breath-held SOTA ([22]) benefits from controlled acquisition and manual cycle labels. By contrast, the proposed pipeline achieves strong performance under more realistic and real-world conditions (free-breathing, automatic segmentation), with top TPR and competitive F1—supporting practical deployability.

### v3

@@ 3.4.3 Multichannel models
The multichannel architecture is shown in Figure 13. The multichannel model includes a support vector machine (SVM) that replaces the hidden layers and output layer of the MLP after training has been complete, leaving only the initial fully connected layer of the MLP. Additionally, for fine-tuning, low-rank adaptation (LoRA) [lora] is utilised for the multichannel model due to this dataset containing more limited data than the other datasets. Both the combination of LoRA for fine-tuning and the SVM help to prevent the model from overfitting to the training set.
@@ 4.1 Single channel models
The PaCMAP [pacmap] plots of an average-performing model from each experiment are shown below to represent how well, on average, those models encode the data before going through the classifier. The data used for these plots is also from the test set, so these models have not been trained on this data. Figure 17, Figure 18, and Figure 19 illustrate the PaCMAP embeddings for the CinC models, with input data sampled at 16kHz with no augmented dataset, 16kHz with augmented dataset and 4.125kHz with augmented dataset, respectively.
@@ 4.2 Multimodal models
Comparing our model that utilises both PCG and ECG signals to other models that were evaluated on training-a, We can see that our model has the greatest performance with the highest for each metric, other than TPR, in Table 16. Although it does not achieve the highest TPR, our proposed model achieves a much higher TNR, than the CNN model from [milan], resulting in all other metrics being greater and hence a better performing model. It is observed that the majority of methods within this comparison utilise the raw signal, as opposed to spectrograms, which is more common for models trained on the entire CinC dataset. However, none of the previous approaches utilised transformer architectures. As training-a is a small dataset, and hence, models trained on it will not generalise well, the results from training on the CinC dataset demonstrate that the method can effectively scale to larger amounts of data while still delivering performance improvements.
@@
CNN [milan] Spectrogram 91.25% 84.17% 98.33% 70.00% 94.40%
RNN [mattone] Raw Signal 91.60% 91.55% 92.00% 91.10% 91.50%
@@ 4.3 Multichannel models
Comparing with the method from the original study [vestpaper], it was observed that their LFCC–SVM approach slightly outperforms the proposed method when evaluated only on the small, breath-held dataset collected from a single round of data acquisition. However, when the original method is re-run on the larger dataset used in this study—which includes an additional round of recordings collected in a different hospital ward—the proposed method achieves superior overall performance. The LFCC–SVM baseline also benefits from cleaner, hand-segmented data recorded while subjects held their breath, whereas the proposed method is trained and evaluated on free-breathing recordings that contain respiratory and background noise, resulting in a noticeable distribution shift relative to the first collection round. As shown in Table 19, the proposed method provides strong improvements in several clinically relevant metrics, most notably achieving the highest TPR (86.47%) and F1-score (82.34%), indicating improved sensitivity to CAD cases and better balance between precision and recall. Although the LFCC–SVM baseline maintains a higher TNR, its performance degrades substantially when extended to the larger dataset, while the proposed method remains stable under noisier real-world conditions. These findings highlight the improved robustness and practicality of the proposed method, which does not require manual heart-cycle segmentation or controlled breath-holding and is therefore better suited for deployment in realistic clinical environments.
@@
LFCC SVM [vestpaper] 80 subjects (40CAD/40NOR), 10s breath held LFCCs 80.44% 80.435% 85.25% 75.62% 81.00%
@@
[…]
@@ 6 Ethics approval and consent
7 Funding
This research did not receive any specific grant from funding agen- cies in the public, commercial, or not-for-profit sectors.